A rectangle is five times as long as it is wide. If the perimeter of the rectangle 240cm, write an equation and solve it to find its length and width.
Answer by texttutoring(324)

Let l=length and w=width
The formula for perimeter is P=2l+2w and we know that l=5w

240=2l+2w

Now replace the l with 5w

240=2(5w)+2w
240=10w+2w
240=12w
w=20

l=5w
l=5*20
l=100

So the width is 20, and the length is 100.